Shipment of 7 television sets contains 3 defective sets. A hotel makes a random purchase
of 3 of the sets. If x is the number of defective sets purchased by the hotel, find the probability
distribution of X. Express the results graphically as a probability histogram.
1
Expert's answer
2021-08-17T17:54:22-0400

The number of possible ways to pick up 3 sets from 7 is "7\cdot6\cdot5=210". Suppose that we want to pick up 1 defective set and two normal. The number of possible choices is: "3\cdot4\cdot3=36", In case we want to collect 2 defective sets and 1 normal, the probability is: "3\cdot2\cdot4=24". In case we want to pick up 3 defective sets, we can make it in "3\cdot2\cdot1=6" ways. In case we want to take 3 normal sets, we can make it in "4\cdot3\cdot2=24" ways. Thus, we get: "P(X=0)=\frac{24}{210}", "P(X=1)=\frac{36}{210}", "P(X=2)=\frac{24}{210}" , "P(X=3)=\frac{6}{210}" .